OpenStudy (anonymous):

f(x)=x^4-5x^3 Curving the Sketch. Find the interval notations for the concavity when it concaves up and concaves down.

OpenStudy (anonymous):

4x^3-15x^2 Critical values: 0,15/4 Increasing (15/4, infinity) Decreasing (-infinity,0]U(0,15/4) Local max: none Local min: 15/4 I found all the way until there and I am stuck.

OpenStudy (anonymous):

Find the second derivative.

OpenStudy (stamp):

Inflection points? \[f'(x)=4x^{3}-15x^{2}\] \[f''(x)=12x^{2}-30x^{1}=6x(2x-5)\] Inflection points f''(x)=0 at x=0 and x=5/2

OpenStudy (anonymous):

\[f(x)=x^4-5x^3 \] \[f'(x)=4x^3-15x^2\] \[f''(x)=12x^2-30x\] At S.P's (Stationary Points) f'(x)=0 \[4x^3-15x^2=0\] \[x^2(4x-15)=0\] \[x=0\] \[y=0\] \[f''(x)=0\] ^^ that point MIGHT be an inflection. Don't take that for granted. \[x=\frac{15}{4}\] \[y=-\frac{16875}{256}\] \[f''(x)>0\] Therefore Minimum T.P (Turning Point) at (\frac{15}{4},-\frac{16875}{256})

OpenStudy (anonymous):

\[(\frac{15}{4},-\frac{16875}{256})\]

OpenStudy (anonymous):

So to find inflection points you would just take the second derivative and make it equal to zero?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Make sure you test the points beside the value so you know how the inflexion point looks like.

OpenStudy (anonymous):

Thank you guys! I got it (: Concave Up: (-Infinity,0)U(5/2,Infinity) Concave down: (0,5/2) Inflection points: x=0, 5/2

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

As regards the point of inflection : you want the third derivative to be non-zero when the second derivative is zero. For instance a line will have second derivative zero everywhere ... but the third derivative will also be zero.

OpenStudy (anonymous):

Thank you for the clarification!